OpenStudy (jmartinez638):

Calculate the following integrals:

OpenStudy (jmartinez638):

\[A) \int\limits_{}^{}(y^2 + 4y - 7)dy\]

OpenStudy (jmartinez638):

\[B) \int\limits_{}^{}\cos(2x)dx\]

OpenStudy (jmartinez638):

\[C) \int\limits_{}^{}\sqrt{x^5}dx\]

OpenStudy (jmartinez638):

D) \[\int\limits_{0}^{\pi}x \sin tdt\]

OpenStudy (anonymous):

A) \[\frac{ y ^{3} }{ 3 }+2y-7y+C\] where C represents a constant B) \[\frac{ 1 }{ 2 }\sin(2x)+C\] where C represents a constant C) \[\frac{ 2 }{ 7 }x ^{\frac{ 7 }{ 2 }}+C\] where C represents a constant D) I think you may have copied this wrong since there should only be one variable in the integrand for single integrals

OpenStudy (jmartinez638):

Yeah, lemme see if I can find the original question(for D).

OpenStudy (jmartinez638):

@Alphabet_sam It turns out that that equation is the "correct" one. Typo perhaps?

OpenStudy (irishboy123):

maybe D is the only interesting one, @Alphabet_Sam ?!?!?

OpenStudy (irishboy123):

for you!!

OpenStudy (jmartinez638):

What do you think @IrishBoy123 ?

OpenStudy (jmartinez638):

Also these here:

OpenStudy (jmartinez638):

\[E) \int\limits_{0}^{x}cosx dx\]

OpenStudy (jmartinez638):

F) \[F) \int\limits_{}^{}\frac{ x-3 }{ \sqrt{x} }dx\]

OpenStudy (jmartinez638):

\[G) \int\limits_{3}^{0}7dt\]

OpenStudy (jmartinez638):

\[H) \int\limits_{-4}^{4}(7t ^{51} - t) dt\]

OpenStudy (jmartinez638):

@jim_thompson5910 I need some help understanding the process of calculating the integrals...

jimthompson5910 (jim_thompson5910):

you need to post these one at a time @jmartinez638

jimthompson5910 (jim_thompson5910):

which one is giving you the most trouble?

OpenStudy (jmartinez638):

I apologize. Well since it's weird, D. H is also giving me a bit of trouble.

jimthompson5910 (jim_thompson5910):

\[\Large \int_{0}^{\pi}x\sin(t)dt\] this?

jimthompson5910 (jim_thompson5910):

I agree with @Alphabet_Sam it's very odd how there are 2 variables here. It suggest there is a typo somewhere

jimthompson5910 (jim_thompson5910):

There's not much we can do really. You'll have to ask your teacher to clear up the problem. I have a feeling they'll say it's a typo too and give you the correct version. Did you want to move onto H now?

OpenStudy (jmartinez638):

Sure!

OpenStudy (irishboy123):

\[\Large \int_{0}^{\pi}x\sin(t)dt\] \[\Large = x \int_{0}^{\pi}\sin(t)dt\]

jimthompson5910 (jim_thompson5910):

x isn't usually a constant, but I guess you could treat it like one

jimthompson5910 (jim_thompson5910):

Part H) \[\Large \int\limits_{-4}^{4}(7t ^{51} - t) dt\] \[\Large \int\limits_{-4}^{4}(7t ^{51})dt - \int\limits_{-4}^{4}(t) dt\] \[\Large 7\int\limits_{-4}^{4}(t ^{51})dt - \int\limits_{-4}^{4}(t) dt\] Do you see how to finish up?

jimthompson5910 (jim_thompson5910):

You'll use this formula \[\Large \int(x^n)dx = \frac{x^{n+1}}{n+1}+C\]

OpenStudy (jmartinez638):

\[7\int\limits_{-4}^{4}(t ^{51})dt - 0\]

jimthompson5910 (jim_thompson5910):

idk how you got that `-0` part

OpenStudy (jmartinez638):

If the first part of the integral minus the second part, the second parts is = 0

OpenStudy (jmartinez638):

maybe i calculated something incorrectly...

jimthompson5910 (jim_thompson5910):

the second part doesn't equal 0

jimthompson5910 (jim_thompson5910):

oh nvm, I miscalculated

jimthompson5910 (jim_thompson5910):

you're right \[\Large \int_{-a}^{a}f(t)dt = 0\] where f(t) is an odd function f(t) = t is an odd function

jimthompson5910 (jim_thompson5910):

so is 7t^(51) since the exponent is odd

OpenStudy (jmartinez638):

So the whole thing is = to zero

jimthompson5910 (jim_thompson5910):

so in reality, you don't even have to find the integral since you can use this shortcut

jimthompson5910 (jim_thompson5910):

yeah

OpenStudy (jmartinez638):

That makes a lot of sense.

OpenStudy (jmartinez638):

I'm going to work on these a little more in a bit. I will open another question if need be, but I doubt there will be any reason.